OpenStudy (133):

I need help if any one can stay with me and help please ASAP!!!! I will give you a medal!!!! @Kfinns99 @Gatorgirl @snowxangel @travisbrown372 @Allsmartpeople!!!!!!!

OpenStudy (133):

Sorry @Kfin99

OpenStudy (133):

Which of the following is equivalent to five and four-ninths divided by seven? seven-ninths one and two-sevenths five-sevenths five and four sixty-thirds

OpenStudy (anonymous):

I believe its one and 2/7

OpenStudy (133):

ok thx and this one:

OpenStudy (133):

Simplify seventeen s plus three t minus nine minus five t plus nine minus twelve s. fourteen s minus two t minus nine three s t five s minus two t fourteen s minus t

OpenStudy (anonymous):

5s-2t

OpenStudy (133):

Thx next one

OpenStudy (133):

Match Term Definition multiply by four, then add three A) y equals four x minus three multiply by three, then subtract four B) y equals four x plus three multiply by four, then subtract three C) y equals three x minus four multiply by three, then add four D) y equals three x plus four

OpenStudy (anonymous):

what class is this? algebra 1?

OpenStudy (133):

Yes

OpenStudy (anonymous):

for connexus?

OpenStudy (133):

no

OpenStudy (133):

Solve three-fourths b minus one-sixth equals one-half for b. b equals eight-ninths b equals one-half b equals one-fourth b equals four-ninths

OpenStudy (133):

@chucky00

OpenStudy (anonymous):

can you reward me? :)

OpenStudy (133):

I did

OpenStudy (133):

Put you as best answer

OpenStudy (anonymous):

thanks! can you help me with my question? or get someone to help? :)

OpenStudy (133):

no I am taking a test and then I have o go into a session so I can't

OpenStudy (133):

*to

OpenStudy (133):

Solve three-fourths b minus one-sixth equals one-half for b. b equals eight-ninths b equals one-half b equals one-fourth b equals four-ninths

OpenStudy (anonymous):

3/4(b) - 1/6 = 1/2 3/4(b) = 1/2 + 1/6 = 3/6 + 1/6 = 4/6 = 2/3 3/4(b) = 2/3 b = 2/3(4/3) = 8/9

OpenStudy (133):

OK thx and the next

OpenStudy (133):

Match Term Definition x divided by negative five is less than 2 A) x is less than negative three negative three x is less than nine B) x is less than negative ten three x is less than negative thirty C) x is greater than negative ten x divided by three is less than negative one D) x is greater than negative three